An engineer claims to have measured the characteristics of a heat engine that takes in 140 J of thermal energy and produces 40 J of useful work.

What is the smallest possible ratio of the temperatures (in kelvin) of the hot and cold reservoirs?

Express your answer using two significant figures. Th/Tc = is the formula used
Answer #1

=1-10 I H

\frac{T_C}{T_H}=1-\frac{W}{Q}=1-\frac{40}{140}=\frac{5}{7}\approx0.71

\frac{\Theta_C+273}{\Theta_H+273}=0.71